OpenStudy (cookiegirl):

Plz help with this question! Estimate the circumference. Use 3.14 or to approximate π. A. 44 cm B. 22 cm C. 88 cm D. 11 cm

OpenStudy (anonymous):

Circumference of a circle is given by: \[\large Circumference = 2 \pi \cdot r\]

OpenStudy (anonymous):

do u hav the circle?

OpenStudy (anonymous):

If you are given with radius, then you are given \(r\) there..

OpenStudy (anonymous):

ya but i don't know what thee radius is...

OpenStudy (anonymous):

the*

OpenStudy (cookiegirl):

@Saisuke<3 i dont have the circle

OpenStudy (anonymous):

But \(\pi\) is related to circle type things..

OpenStudy (anonymous):

mkay then do u hav the radius or the diameter?

OpenStudy (cookiegirl):

the radius is 7cm

OpenStudy (anonymous):

ok so if the radius is 7 then do u know the diameter?

OpenStudy (anonymous):

See, you are given \(r\) here..

OpenStudy (cookiegirl):

no it doesn't show it

OpenStudy (anonymous):



OpenStudy (cookiegirl):

it looks somewhat like that

OpenStudy (anonymous):

and u need to find circumference

OpenStudy (anonymous):

so do u know what the formula for circumference is

OpenStudy (cookiegirl):

it doesn't say...... this is so confusing :(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its 2\(\pi\) r

OpenStudy (anonymous):

i can help!!

OpenStudy (anonymous):

so basically, you do 2 times \(\pi\) times the radius

OpenStudy (anonymous):

ok so whats the radious?

OpenStudy (anonymous):

so first, you might want to do 2 times \(\pi\)

OpenStudy (anonymous):

do you know what that is?

OpenStudy (cookiegirl):

sort of...

OpenStudy (anonymous):

ok so do the math and tell me what u get

OpenStudy (anonymous):

2•\(\pi\) = ?

OpenStudy (anonymous):